# It’s complicated. The relationship of algorithmic fairness and non-discrimination regulations in the EU AI Act

### v2

@@ Abstract.
What constitutes a fair decision? This question is not only difficult for humans but becomes more challenging when Artificial Intelligence (AI) models are used. In light of discriminatory algorithmic behaviors, the EU has recently passed the AI Act, which mandates specific rules for AI models, incorporating both traditional legal non-discrimination regulations and machine learning based algorithmic fairness concepts. This paper aims to bridge these two different concepts in the AI Act through: First a high-level introduction of both concepts targeting legal and computer science-oriented scholars, and second an in-depth analysis of the AI Act’s relationship between legal non-discrimination regulations and algorithmic fairness. Our analysis reveals three key findings: (1.), most non-discrimination regulations target only high-risk AI systems. (2.), the regulation of high-risk systems encompasses both data input requirements and output monitoring, though these regulations are partly inconsistent and raise questions of computational feasibility. (3.) Regulations for General Purpose AI Models, such as Large Language Models that are not simultaneously classified as high-risk systems, currently lack specificity compared to other regulations. Based on these findings, we recommend developing more specific auditing and testing methodologies for AI systems. This paper aims to serve as a foundation for future interdisciplinary collaboration between legal scholars and computer science-oriented machine learning researchers studying discrimination in AI systems.
EU AI Act, Algorithmic Fairness, Discrimination, Bias, Fundamental Rights
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 2025 ACM Conference on Fairness, Accountability, and Transparency; ;
1. Introduction
@@
How can we ensure fair algorithms in the context of AI Systems? When regulating algorithms and, more specifically, Artificial Intelligence (AI) systems, this question becomes fundamental. While the question is also crucial in an analog world, it becomes increasingly pressing in the digital world. The European Union (EU) has become one of the forerunners of regulations in the digital age with regulatory frameworks, such as the Digital Service Act [^1] Regulation (EU) 2022/2065., the Digital Markets Act [^2] Regulation (EU) 2022/1925., or the Data Act [^3] Regulation (EU) 2023/2854.. In order to maintain fair data processing in the machine learning domain, the European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^4] Regulation (EU) 2024/1689.. The aim of this AI regulation is to maintain a level playing field for “ethical principles” (Nolte et al., 2024) within and outside the European Union (see recital 27 AIA).
@@
Fair algorithmic processing matters also from a computational perspective. Over the past decade, algorithmic fairness has become a well-established field within machine learning that focuses on defining, mitigating, and evaluating discriminatory behavior of Artificial Intelligence models (Pessach and Shmueli, 2023). The importance becomes clear when looking at several discriminatory behaviors of algorithms in the past, ranging from hiring (Dastin, 2022) to Social Welfare systems (Hadwick and Lan, 2021). In recent years, large generative (multimodal) models, particularly Large Language Models (LLMs) with their high accessibility and wide range of applications have posed significant new challenges to the algorithmic fairness domain (Chu et al., 2024; Kotek et al., 2023).
@@
The relationship between computer science and legislation for efficient regulations within the AI Act is complex. While the interplay between algorithmic fairness and non-discrimination law is often unclear (Weerts et al., 2023), efficient AI legislation requires interdisciplinary collaboration connecting the legal domain and computer science. Although, legal and technical concerns regarding algorithmic fairness and non-discrimination law have a long history (Barocas and Selbst, 2016; Hacker, 2018; Wachter et al., 2021), recent developments have accelerated the pace. An increasing number of companies and laypersons are using these technologies, and with increased usage, the likelihood of misuse also rises. Thus, the need for efficient and effective regulations to protect individuals is substantial.
@@
Previous literature: Many papers on discrimination and AI, limited amount on the interaction of the AI Act and discrimination. Discrimination in algorithms is not a novel phenomenon and was analyzed in a pre-AI Act era. Already in the 1980s, algorithmic discrimination was observed in admission settings (Connors et al., 1981; Williams et al., 1981), and gender inequalities in educational software were discussed (Huff and Cooper, 1987). In 1996, researchers pointed out the levels at which such technical constraints and social institutions’ biases can occur (Friedman and Nissenbaum, 1996). More recently, the interaction between AI and automated decision-making systems with traditional European and US non-discrimination law has been studied extensively. Barocas and Selbst (2016) presented a taxonomy of different sources of discrimination and their impact on humans. The relationship among various European non-discrimination laws has also been studied extensively (Wachter et al., 2021, 2020; Hacker, 2018; Weerts et al., 2023; Xenidis and Senden, 2020). For example, the seminal study by Wachter et al. (2021) presented a fairness metric that connects algorithmic fairness to the jurisprudence of the European Court of Justice in relation to the General Data Protection Regulation. Similarly, the connection between non-discrimination law and the GDPR has been analyzed to “unlock the algorithmic black box” (Hacker, 2018). Naturally, since these papers were published before the AI Act was passed, they do not contain references to it.
@@
Recently, several papers have already addressed the AI Act. Short(er) studies such as (Deck et al., 2024a; Ruohonen, 2024) have analyzed the relationship between fairness and the AI Act. Additionally, sections within papers of broader scope beyond pure non-discrimination analyses have mentioned aspects of the relationship between algorithmic fairness and the AI Act (Wachter, 2024; Hacker et al., 2024b; Novelli et al., 2024). The AI Act and its relation to gender equality and non-discrimination law have also been discussed (Lütz, 2024). Most similar to our paper in terms of the AI Act, Bosoer et al. (2023) investigated the non-discrimination regulations in the draft of the AI Act without a strong focus on the interaction with computer science. However, their focus was on the draft, whereas our paper focuses on the substantially different final version. Most similar in terms of the interaction between computer science and legal research is the paper of Weerts et al. (2023). In contrast to our work they did not focus on the AI Act.
@@
Our contributions: From regulatory clarity to vagueness This paper aims to foster and extend an interdisciplinary view on explicit and implicit fairness regulations in the AI Act regarding algorithmic fairness in the ML domain. We aim to introduce the field of algorithmic fairness to legally oriented researchers in Section 2.1. Afterwards, we introduce non-discrimination law to computer scientists in Section 2.2 and discuss the truncated relationship between the two fields in Section 2.3.
For the in-depth analysis of the EU AI Act non-discrimination regulations in Section 3, our contributions are as follows (Figure 1 shows our results and findings in a nutshell):
  1. (1)
    We present the history, scope, and intentions of non-discrimination regulations in the AI Act (Sec. 3.1).
  2. (2)
    We identify which issues of algorithmic fairness are addressed explicitly and implicitly within the AI Act (Sec. 3.2).
  3. (3)
@@
    We find that high-risk AI Systems (Article 6 AIA) are most extensively regulated regarding non-discrimination. Regulations are partly explicit and partly implicit. Overall the regulations are inconsistent with respect to algorithmic fairness, with regulations on input and output data at different stages (Sec. 3.3).
  4. (4)
    Non-discrimination regulations of General Purpose AI models (GPAI) in Article 53-56 AIA remain insufficiently specific and lack clear definition (Sec. 3.4).
@@
We summarize our findings and give recommendations for future research in Section 4. Additionally, based on our results, we propose approaches to overcome some of the challenges and sketch possible actions to address the inconsistencies. This includes a call for closer interactions between computer scientists and legal researchers.
Figure 1. Results of our analysis of the EU AI Act’s non-discrimination regulation and algorithmic fairness in a nutshell: Low/medium risk systems [^5] The AI Act also includes a third category of AI Models in Article 51 AIA: certain AI Systems. However, certain AI Systems are not target of any specific non-discrimination regulation and thus not included in our analysis.  are not regulated in terms of non-discrimination obligations within the AI Act. Key terms for high-risk systems are the partly well-defined concept of non-discrimination and bias (less well defined). High-risk systems, as well as GPAI models, are also regulated with respect to Fundamental Rights. GPAI models, with the key term systemic risk, show the lowest regulatory clarity. Overall, regulatory clarity for non-discrimination matters decreases from high-risk systems to GPAI models.
2. A primer on non-discrimination from an AI and a legal perspective
@@
However, the complexity in interdisciplinary work extends beyond pure knowledge. The challenges begin with terminology: For instance, in computer science, “fairness” is a term that can refer to different desiderata that aim to prevent socially or morally undesirable behavior or outcomes of algorithms. Although fairness has been discussed as principle in economic law contexts (Scheuerer, 2023), it is not a specific legal term. The arguably closest legal counterpart term is “non-discrimination”, which focuses on preventing unfair treatment based on characteristics such as race, gender, or other attributes on legal grounds. The interaction between both key concepts from law and computer science is demanding.
@@
Thus, we highlight the differences and common grounds of algorithmic fairness and non-discrimination regulations in the following sections. After presenting an introduction to algorithmic fairness in Section 2.1 — for legal researchers — and an introduction to EU non-discrimination law in Section 2.2 — for computer science oriented researchers —, we try to bridge the gap between these two areas in Section 2.3.
2.1. Algorithmic fairness: A gentle introduction
@@
Algorithmic Fairness has emerged as an established computer science and AI related field over the past years. Algorithmic fairness focuses on uncovering and rectifying disadvantageous treatment of individuals or groups in machine learning models (Mitchell et al., 2021). This treatment must be considered within appropriate social, theoretical, and legal contexts (Pessach and Shmueli, 2023) and includes evaluation and auditing methods to test for unfairness.
@@
Origins of unfairness in algorithms. The origins of unfairness are manifold. Unfair algorithmic behavior can have different sources. In order to discuss the origins, another — not well defined — term from computer science is often used: biases. There exist many different bias definitions. Since the EU AI Act does not define it itself, we use a definition from the EU commission from 2021. According to this, a bias can be defined as “[…] bias describes systematic and repeatable errors in a computer system that create unfair outcomes, such as favouring one arbitrary group of users over others” (Commission, 2021). Biases are primarily discussed within data used for AI models and can be sorted into different categories. Most researchers emphasize the different aspects of data inequalities when discussing algorithmic fairness and sort data biases into different categories, such as the over-representation of specific groups in datasets (Zou and Schiebinger, 2018). Hacker (2018) sorts the biases in data into two groups: biased training data vs. unequal ground truth; others use more categories. Barocas and Selbst (2016) uses five categories to map biases, while Mehrabi et al. (2021) uses 21 different categories for biases. It is important to note that the data are only — one albeit important — source of biases. For example, design decisions in the algorithms or structural power asymmetries also lead to algorithmic unfairness (Mehrabi et al., 2021; Gebru et al., 2021; Sousa e Silva, 2024).
@@
Approaches to algorithmic fairness: Quantification. Based on the different understandings of unfairness and for the purpose of uncovering and rectifying unjustified treatment between groups in algorithms, so-called fairness metrics have been proposed (Corbett-Davies et al., 2023; Castelnovo et al., 2022; Verma and Rubin, 2018). The core idea of fairness metrics is to quantify algorithmic outputs and make them comparable through numerical measurements. Most metrics try to assess whether an output of an ML system is unfair to individuals (Dwork et al., 2012) or groups of people (Barocas et al., 2023; Binns, 2020). They provide a quantitative measure intended to indicate whether an algorithm shows unjustified unequal treatment. Fairness metrics also include moral norms (Deck et al., 2024a; Hellman, 2020).
@@
The emergence of the field of algorithmic fairness has challenged existing approaches to ML by highlighting the need to contextualize them, decide on and reason for the optimization criteria chosen. First, it has been shown that an apparent objective number needs to be contextualized to the application under consideration (Wachter et al., 2021). Furthermore, some metrics are, from a mathematical viewpoint incompatible with each other [^6] Please note that they are also incompatible from a moral point of view (Heidari et al., 2019)., thus some fairness metrics cannot be fulfilled at the same time (Chouldechova, 2017; Kleinberg et al., 2016). This can lead to an effect of “d-hacking” or “fairness hacking” where users can choose their favorite metric to create the (technical and mathematical) impression that the algorithm is fair (Black et al., 2024; Meding and Hagendorff, 2024). Fairness metrics for mitigation can play a role at different stages of the development, testing and deployment of an algorithm. The pre-processing stage, the in-training stage, and the post-processing phases are differentiated (Barocas et al., 2023; Binns, 2018). In the pre-processing phase, the input data itself is altered to ensure fair processing (Caton and Haas, 2024; Kamiran and Calders, 2012). For the in-training phase, the optimization process during the training of an ML model is adjusted to include fairness as an optimization goal. Finally, in the post-processing phase, the output of an already trained model is adapted, see Figure 2 in the appendix. The first and the last approach makes it possible to perform fairness analyzes even if one only has black box access to the model (Caton and Haas, 2024).
@@
The era of LLMs further complicates the situation. LLMs have the advantage — or disadvantage depending on the viewpoint — that their exact use case is most of the time not predefined. They are trained on large amounts of various data. This diversity makes them later applicable to different contexts, from providing cookie recipes to coding exercises. However, the diversity then also includes variations in unfairness. Algorithmic fairness in relation to LLMs is thus somewhat different from algorithmic fairness in classical AI (Doan et al., 2024; Chu et al., 2024). Some researchers have applied classical fairness metrics in classification settings to LLMs (Chhikara et al., 2024). Additionally, it was shown that LLM-specific issues arise due to the use of massive data and processing (Kotek et al., 2023; Ferrara, 2023; Navigli et al., 2023; Huang et al., 2024) at various steps of algorithmic development. Thus, it was argued that LLMs cannot yield fair outcomes at all (Anthis et al., 2024).
2.2. EU non-discrimination Law: A gentle introduction, too.
Law is different from computer science. Nevertheless — of course — legal researchers have also discussed different concepts of discrimination. In contrast to the term “fairness”, the term “(non-)discrimination” is a specific legal concept  [^7] For an in-depth introduction we refer to the paper of Weerts et al. (2023) or Wachter et al. (2020).. For a better understanding of these legal concepts, we first describe traditional EU non-discrimination law as well as the Fundamental Rights approach of the EU. Both influence the AI Act.
Direct and indirect discrimination in EU Law. EU Law has a long history of non-discrimination regulations and many different laws include non-discrimination regulations [^8] See for an overview: https://commission.europa.eu/aid-development-cooperation-fundamental-rights/your-fundamental-rights-eu/know-your-rights/equality/non-discrimination_en. Primary EU law already includes non-discrimination regulations, one example is the prohibition of unequal payment based on gender (Article 157(1) TFEU). In order to legally assess discrimination under EU Law, traditionally two different versions of discrimination are differentiated: direct and indirect discrimination (Zuiderveen Borgesius et al., 2024; Wachter et al., 2020). In context of the US law, these two concepts correspond to disparate treatment (direct discrimination) and disparate impact (indirect discrimination) (Weerts et al., 2023). However, importantly, EU-based direct discrimination does not need — compared to US-based disparate treatment — any intentional wrongdoing (Weerts et al., 2023; Xenidis and Senden, 2020; Adams-Prassl et al., 2023).
The AI Act is based on the legal approach of discrimination-related regulations (Deck et al., 2024a). Thus, when we analyze the EU AI Act regulations in this paper, we will minimize the usage of the vague CS related concept of fairness and instead focus on the legally related term of non-discrimination.
Direct discrimination is defined as situations in which “one person is treated less favourably than another is, has been or would be treated in a comparable situation” (Article 2(2)(a) Directive 2000/43/EC). This means that individuals are discriminated against by using a (listed) protected attribute. In algorithmic contexts, an example would be a simple algorithm that grants anyone below 50 years a loan and denies it to anyone above 50 years. In this setting, only the protected attribute age is used to define the loan threshold. People would be directly discriminated against based on their age.
Indirect discrimination is much more challenging to tackle. In indirect discrimination cases, seemingly neutral attributes are used, but they rely on a protected attribute. An example of indirect discrimination would be if not the age is used but the amount of gray hair, which is usually highly correlated with age. If an algorithm does not use age directly but the amount of gray hair, then people are indirectly discriminated against based on their age. For indirect discrimination, it is important to note that it can be justified through legitimate aims and appropriate means (Zuiderveen Borgesius et al., 2024).
Also, a third category of discrimination needs to be taken into account [^9] Please note that other non-discrimination rules in sector-specific regulations, such as the EU Consumer Credit Directive, are beyond the scope of our paper., which challenges traditional non-discrimination law: intersectional discrimination. Intersectional discrimination concerns cases “originating in several inextricably linked vectors of disadvantage” (Xenidis, 2023). This becomes especially important since (modern) AI models do not use single variables as input — such as our simplified loan example based on the single variable above — but instead use many different aspects as input. This can lead to effects where discrimination may occur at the intersection of gender and age. Importantly, if you only investigate age and gender separately, no discrimination would be recognized (Weerts et al., 2023). It is an open question if and how, at the current stage, the European Court of Justice recognizes intersectional discrimination as a form of discrimination (Weerts et al., 2023; Xenidis, 2023; Atrey, 2018).
With these three types of discrimination based on non-discrimination law, the last concept that needs to be discussed is the concept of Fundamental Rights which is mentioned in multiple articles in the AI Act.
@@
Fundamental Rights in EU law. The Charter of Fundamental Rights of the European Union (CFR) is the main Fundamental Rights regulation within EU law . Through Article 6(1) of the Treaty on European Union, the Charter is part of the primary law of the EU.
@@
The CFR also includes non-discrimination law. For non-discrimination setting obligations, Article 21 CFR states: “Any discrimination based on any ground such as sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age or sexual orientation shall be prohibited.”

### v3

# It’s complicated. The relationship of algorithmic fairness and non-discrimination regulations for high-risk systems in the EU AI Act

@@ Abstract
What constitutes a fair decision? This question is not only difficult for humans but becomes more challenging when Artificial Intelligence (AI) models are used. In light of discriminatory algorithmic behaviors, the EU has recently passed the AI Act, which mandates specific rules for high-risk systems, incorporating both traditional legal non-discrimination regulations and machine learning based algorithmic fairness concepts. This paper aims to bridge these two different concepts in the AI Act through: First, a necessary high-level introduction of both concepts targeting legal and computer science-oriented scholars, and second, an in-depth analysis of the AI Act’s relationship between legal non-discrimination regulations and algorithmic fairness. Our analysis reveals three key findings: (1.) Most non-discrimination regulations target only high-risk AI systems. (2.) The regulation of high-risk systems encompasses both data input requirements and output monitoring, though these regulations are partly inconsistent and raise questions of computational feasibility. (3.) Finally, we consider the possible (future) interaction of classical EU non-discrimination law and the AI Act regulations. We recommend developing more specific auditing and testing methodologies for AI systems. This paper aims to serve as a foundation for future interdisciplinary collaboration between legal scholars and computer science-oriented machine learning researchers studying discrimination in AI systems.
1 Introduction
@@
How can we ensure fair algorithms in the context of AI systems? When regulating algorithms and, more specifically, Artificial Intelligence (AI) systems, this question becomes fundamental. While the question is also crucial in an analog world, it becomes increasingly pressing in the digital world. The European Union (EU) has become one of the forerunners of regulations in the digital age with regulatory frameworks, such as the Digital Services Act [^1] Regulation (EU) 2022/2065., the Digital Markets Act [^2] Regulation (EU) 2022/1925., or the Data Act [^3] Regulation (EU) 2023/2854.. In order to maintain fair data processing in the machine learning domain, the European Union (EU) recently adopted the Artificial Intelligence Act (AIA) [^4] Regulation (EU) 2024/1689.. The aim of this AI regulation is to maintain a level playing field for “ethical principles” (Nolte, Rateike, and Finck 2024) within and outside the European Union (see recital 27 AIA).
@@
Fair algorithmic processing also matters from a computational perspective. Over the past decade, algorithmic fairness has become a well-established field within machine learning that focuses on defining, mitigating, and evaluating discriminatory behavior of Artificial Intelligence models (Pessach and Shmueli 2023). The importance becomes evident when looking at several discriminatory behaviors of algorithms in the past, ranging from hiring (Dastin 2022) to social welfare systems (Hadwick and Lan 2021). In recent years, large generative (multimodal) models, particularly Large Language Models (LLMs) with their high accessibility and wide range of applications have posed significant new challenges to the algorithmic fairness domain (Chu, Wang, and Zhang 2024; Kotek, Dockum, and Sun 2023).
@@
The relationship between computer science and legislation within the AI Act is complex. While the interplay between algorithmic fairness and non-discrimination law is often unclear (Weerts et al. 2023), efficient AI legislation requires interdisciplinary collaboration connecting the legal domain and computer science. Although legal and technical concerns regarding algorithmic fairness and non-discrimination law have existed for a long time (Barocas and Selbst 2016; Hacker 2018; Wachter, Mittelstadt, and Russell 2021), recent developments have accelerated the pace. An increasing number of companies and laypersons are using these technologies, and with increased usage, the likelihood of misuse also rises. Thus, the need for efficient and effective regulations to protect individuals is substantial.
@@
Previous literature. Discrimination through algorithms is not a novel phenomenon and was analyzed in a pre-AI Act era. Already in the 1980s, algorithmic discrimination was observed in admissions settings (Connors et al. 1981; Williams, Werth, and Wolff 1981), and gender inequalities in educational software were discussed (Huff and Cooper 1987). In 1996, researchers pointed out the levels at which such technical constraints and social institutions’ biases can occur (Friedman and Nissenbaum 1996). More recently, the interaction between AI and automated decision-making systems with traditional European and US non-discrimination law has been studied extensively. Barocas and Selbst (2016) presented a taxonomy of different sources of discrimination and their impact on humans. The relationship among various European non-discrimination laws has also been studied (Wachter, Mittelstadt, and Russell 2021, 2020; Hacker 2018; Weerts et al. 2023; Xenidis and Senden 2020; Lewis et al. 2025). For example, the seminal study by Wachter, Mittelstadt, and Russell (2021) presented a fairness metric that connects algorithmic fairness to the jurisprudence of the European Court of Justice in relation to the General Data Protection Regulation. Similarly, the connection between non-discrimination law and the GDPR has been analyzed to “unlock the algorithmic black box” (Hacker 2018). Naturally, since these papers were published before the AI Act was passed, they do not contain references to it.
@@
Recently, several papers have already addressed the AI Act. Short studies such as Deck et al. (2024a); Ruohonen (2024) have analyzed the relationship between fairness and the AI Act. Additionally, sections within papers of broader scope beyond pure non-discrimination analyses have mentioned aspects of the relationship between algorithmic fairness and the AI Act (Wachter 2024; Hacker et al. 2024; Novelli et al. 2024). The AI Act and its relation to gender equality and non-discrimination law have also been discussed (Lütz 2024). Among these, the work by Bosoer, Cantero Gamito, and Rubio-Marin (2023) is most similar to ours in terms of its focus on the AI Act; however, it investigates non-discrimination regulations in the draft of the AI Act without a strong focus on the interaction with computer science. Most similar in terms of the interaction between computer science and legal research is the paper of Weerts et al. (2023). In contrast to our work, they did not focus on the AI Act.
@@
Our contributions. This paper aims to foster and extend an interdisciplinary view on explicit and implicit fairness regulations in the AI Act regarding algorithmic fairness in the ML domain. We aim to introduce the field of algorithmic fairness to legally oriented researchers and discuss the fragmented relationship between the two fields. These sections are necessary for a broader audience to follow our reasoning.
For the in-depth analysis of the EU AI Act’s non-discrimination regulations in Section 3, our contributions are as follows
  1. 1.
    We present the history, scope, and intentions of non-discrimination regulations in the AI Act
  2. 2.
@@
    We find that high-risk AI systems (Article 6 AIA) are most thoroughly regulated regarding non-discrimination. We discuss the relation between high-risk systems’ regulations and algorithmic fairness. We find that the AI Act mainly regulates input data compared to the output data of AI systems.
  3. 3.
    We discuss challenges for classical EU non-discrimination at the intersection of the AI Act, since classical EU non-discrimination law will still be important in AI contexts.
@@
In the last section, we summarize our findings and give recommendations for future research. Additionally, based on our results, we suggest approaches to overcome some challenges and propose possible actions to address them.
2 A primer on non-discrimination from an AI and a legal perspective
@@
However, the complexity in interdisciplinary work extends beyond pure knowledge. The challenges begin with terminology: For instance, in computer science, “fairness” is a term that can refer to different desiderata that aim to prevent socially or morally undesirable behavior or outcomes of algorithms. Although fairness has been discussed as a principle in economic law contexts (Scheuerer 2023), it is not a specific legal term. The arguably closest legal counterpart term is “non-discrimination”, which focuses on preventing unfair treatment based on characteristics such as race, gender, or other attributes on legal grounds. The interaction between these key concepts from law and computer science is demanding.
@@
Thus, we highlight the differences and common grounds of algorithmic fairness and non-discrimination regulations in the following sections. After presenting an introduction to algorithmic fairness — for legal researchers — and an introduction to EU non-discrimination law — for computer science oriented researchers —, we attempt to bridge the gap between these two areas.
Algorithmic fairness: A gentle introduction
@@
Algorithmic fairness has emerged as an established computer science and AI-related field over the past years. Algorithmic fairness focuses on uncovering and rectifying disadvantageous treatment of individuals or groups in machine learning models (Mitchell et al. 2021). This treatment must be considered within appropriate social, theoretical, and legal contexts (Pessach and Shmueli 2023) and includes evaluation and auditing methods to test for unfairness.
@@
Origins of unfairness in algorithms. Unfair algorithmic behavior can have different sources. In order to discuss the origins, another — not well-defined — a term from computer science is often used: bias. There exist many different bias definitions. Since the EU AI Act does not define it itself, we use a definition from the EU Commission from 2021. According to this, a bias can be defined as “[…] bias describes systematic and repeatable errors in a computer system that create unfair outcomes, such as favouring one arbitrary group of users over others” (Commission 2021). Biases are primarily discussed within the data used for AI models and can be sorted into different categories. Most researchers emphasize the different aspects of data inequalities when discussing algorithmic fairness and sort data biases into different categories, such as the over-representation of specific groups in datasets (Zou and Schiebinger 2018). Hacker (2018) categorizes the biases in data into two groups: biased training data versus unequal ground truth; other authors use more categories. Barocas and Selbst (2016) uses five categories to map biases, while Mehrabi et al. (2021) uses 21 different categories for biases. It is important to note that the data are only one — albeit important — source of biases. For example, design decisions in the algorithms or structural power asymmetries also lead to algorithmic unfairness (Mehrabi et al. 2021; Gebru et al. 2021; Sousa e Silva 2024).
@@
Approaches to algorithmic fairness: Quantification. Based on the different understandings of unfairness and for the purpose of uncovering and rectifying unjustified treatment between groups in algorithms, so-called fairness metrics have been proposed (Corbett-Davies et al. 2023; Castelnovo et al. 2022; Verma and Rubin 2018). The core idea of fairness metrics is to quantify algorithmic outputs and make them comparable through numerical measurements. Most metrics try to assess whether an output of an ML system is unfair to individuals (Dwork et al. 2012) or groups of people (Barocas, Hardt, and Narayanan 2023; Binns 2020). They provide a quantitative measure intended to indicate whether an algorithm demonstrates unjustified unequal treatment. Fairness metrics also include moral norms (Deck et al. 2024a; Hellman 2020).
@@
The emergence of the field of algorithmic fairness has challenged existing approaches to ML by highlighting the need to contextualize them, decide on, and provide a rationale for the optimization criteria chosen. First, it has been shown that an apparently objective number needs to be contextualized to the application under consideration (Wachter, Mittelstadt, and Russell 2021). Furthermore, some metrics are, from a mathematical viewpoint incompatible with each other [^5] They are also incompatible from a moral point of view (Heidari et al. 2019)., thus these fairness metrics cannot be fulfilled at the same time (Chouldechova 2017; Kleinberg, Mullainathan, and Raghavan 2016). This can lead to an effect of “d-hacking” or “fairness hacking” where users can choose their favorite metric to create the (technical and mathematical) impression that the algorithm is fair (Black, Gillis, and Hall 2024; Meding and Hagendorff 2024). Fairness metrics for mitigation can play a role at different stages of the development, testing and deployment of an algorithm. The pre-processing stage, the in-training stage, and the post-processing phases are differentiated (Barocas, Hardt, and Narayanan 2023; Binns 2018). In the pre-processing phase, the input data itself is altered to ensure fair processing (Caton and Haas 2024; Kamiran and Calders 2012). For the in-training phase, the optimization process during the training of an ML model is adjusted to include fairness as an optimization goal. Finally, in the post-processing phase, the output of an already trained model is adapted. The first and the last approaches make it possible to perform fairness analyses even if one only has black box access to the model (Caton and Haas 2024).
@@
The era of LLMs raises new questions. LLMs have the advantage — or disadvantage, depending on the viewpoint — that their exact use case is most of the time not predefined, and they are trained on large amounts of various data. This diversity makes them later applicable to different contexts, from providing cookie recipes to coding exercises. However, this diversity also includes variations in unfairness. Algorithmic fairness in relation to LLMs is thus somewhat different from algorithmic fairness in classical AI (Doan et al. 2024; Chu, Wang, and Zhang 2024). Some researchers have applied classical fairness metrics in classification settings to LLMs (Chhikara et al. 2024). Additionally, it has been shown that LLM-specific issues arise due to the use of massive data and processing (Kotek, Dockum, and Sun 2023; Ferrara 2023; Navigli, Conia, and Ross 2023; Huang et al. 2024) at various steps of algorithmic development. Thus, it was argued that LLMs cannot yield fair outcomes at all (Anthis et al. 2024).
EU non-discrimination Law: A gentle introduction
We first describe fundamental rights as well as traditional EU non-discrimination law. Both influence the AI Act.
@@
Fundamental rights in EU law. The Charter of Fundamental Rights of the European Union (CFR) is the main fundamental rights regulation within EU law [^6] The CFR should not be confused with European Convention on Human Rights (ECHR). The ECHR was not developed by the EU. Nevertheless, the EU needs to accede to the ECHR. The exact relationship of the ECHR to EU law is complex, see (Brittain 2015) for further discussion.. Through Article 6(1) of the Treaty on European Union, the Charter is part of the primary law of the EU.
@@
The CFR includes non-discrimination law. For non-discrimination setting obligations, Article 21 CFR states: “Any discrimination based on any ground such as sex, race, color, ethnic or social origin, genetic features, language, religion or belief, political or any other opinion, membership of a national minority, property, birth, disability, age or sexual orientation shall be prohibited.” Thus, only discrimination by these listed attributes is targeted.